### v2

@@ 2.1 Data
The BraTS-PEDs dataset includes a retrospective multi-institutional cohort of conventional/structural magnetic resonance imaging (MRI) sequences, including pre- and post-gadolinium T1-weighted (labeled as T1 and T1CE), T2-weighted (T2), and T2-weighted fluid attenuated inversion recovery (T2-FLAIR) images, from 228 pediatric high-grade glioma. These conventional multiparametric MRI (mpMRI) sequences are commonly acquired as part of standard clinical imaging for brain tumors. However, the image acquisition protocols and MRI equipment differ across different institutions, resulting in heterogeneity in image quality in the provided cohort. Inclusion criteria comprised of pediatric subjects with: (1) histologically-approved high-grade glioma, i.e., high-grade astrocytoma and diffuse midline glioma (DMG), including radiologically or histologically-proven diffuse intrinsic pontine glioma (DIPG); (2) availability of all four structural mpMRI sequences on treatment-naive imaging sessions. Exclusion criteria consisted of: (1) images assessed to be of low quality or with artifacts that would not allow for reliable tumor segmentation; and (2) infants younger than one month of age. Data for n = 228 patients was obtained through CBTN (n = 122), DMG/ DIPG Registry (n = 256), Boston’s Children Hospital (n = 61), and Yale University (n = 27).
@@ 2.1.1 Imaging Data Description
    Enhancing tumor (ET), described by areas with enhancement (brightness) on T1 post-contrast images as compared to T1 pre-contrast. In case of mild enhancement, checking the signal intensity of normal brain structure can be helpful.
@@
    Nonenhancing tumor (NET), defined as any other abnormal signal intensity within the tumorous region that cannot be defined as enhancing or cystic. For example, the abnormal signal intensity on T1, T2-FLAIR, and T2 that is not enhancing on T1CE should be considered as nonenhancing portion.
@@
    Cystic component (CC), typically appearing with hyperintense signal (very bright) on T2 and hypointense signal (dark) on T1CE. The cystic portion should be within the tumor, either centrally or peripherally (as compared to ED which is peritumoral). The brightness of CC is here defined as comparable or close to cerebrospinal fluid (CSF).
@@
    Peritumoral edema (ED), defined by the abnormal hyperintense signal (very bright) on FLAIR scans. ED is finger-like spreading that preserves underlying brain structure and surrounds the tumor.
@@ 2.2 Performance Evaluation
As described earlier and exemplified in Figure 2 (bottom panel), for the BraTS-PEDs 2023 challenge, the regions to evaluate the performances are: i) the “enhancing tumor” (ET), ii) the “enhancing tumor/cystic component/necrosis” (NC) and iii) the entire tumorous region, or the so-called “whole tumor” (WT).
@@
    Online evaluation platform (Synapse)

### v4

@@ 2.1 Data
The BraTS-PEDs dataset includes a retrospective multi-institutional cohort of conventional/structural magnetic resonance imaging (MRI) sequences, including pre- and post-gadolinium T1-weighted (labeled as T1 and T1CE), T2-weighted (T2), and T2-weighted fluid attenuated inversion recovery (T2-FLAIR) images, from 464 pediatric high-grade glioma. These conventional multiparametric MRI (mpMRI) sequences are commonly acquired as part of standard clinical imaging for brain tumors. However, the image acquisition protocols and MRI equipment differ across different institutions, resulting in heterogeneity in image quality in the provided cohort. Inclusion criteria comprised of pediatric subjects with: (1) histologically-approved high-grade glioma, i.e., high-grade astrocytoma and diffuse midline glioma (DMG), including radiologically or histologically-proven diffuse intrinsic pontine glioma (DIPG); (2) availability of all four structural mpMRI sequences on treatment-naive imaging sessions. Exclusion criteria consisted of: (1) images assessed to be of low quality or with artifacts that would not allow for reliable tumor segmentation; and (2) infants younger than one month of age. Data for 464 patients was obtained through CBTN (n = 120), DMG/ DIPG Registry (n = 256), Boston’s Children Hospital (n = 61), and Yale University (n = 27).
@@ 2.1.1 Imaging Data Description
    Enhancing tumor (ET; label 1; value = 1), described by areas with enhancement (brightness) on T1 post-contrast images as compared to T1 pre-contrast. In case of mild enhancement, checking the signal intensity of normal brain structure can be helpful.
@@
    Nonenhancing tumor (NET; label 2; value = 2), defined as any other abnormal signal intensity within the tumorous region that cannot be defined as enhancing or cystic. For example, the abnormal signal intensity on T1, T2-FLAIR, and T2 that is not enhancing on T1CE should be considered as nonenhancing portion.
@@
    Cystic component (CC; label 3; value = 3), typically appearing with hyperintense signal (very bright) on T2 and hypointense signal (dark) on T1CE. The cystic portion should be within the tumor, either centrally or peripherally (as compared to ED which is peritumoral). The brightness of CC is here defined as comparable or close to cerebrospinal fluid (CSF).
@@
    Peritumoral edema (ED; label 4; value = 4), defined by the abnormal hyperintense signal (very bright) on FLAIR scans. ED is finger-like spreading that preserves underlying brain structure and surrounds the tumor.
    The BraTS-PEDs 2024 can be accessed at https://www.synapse.org/Synapse:syn58894466.
Important Note: The BraTS-PEDs 2024 is not skull-stripped. If you need to skull strip the images, you may use our open-access model provided at one of the following repositories:
  1. 1.
  2. 2.
    We caution the participants about using skull-stripping in training their models, as the validation and testing data are only defaced and not skull-stripped.
@@ 2.2 Performance Evaluation
Important Note: The BraTS-PEDs 2024 evaluation is different from the 2023 challenge.
For the BraTS-PEDs 2024 challenge, the regions to evaluate the performances are the: i) “enhancing tumor” (ET), ii) “non-enhancing tumor” (NET), iii) “cystic component” (CC), iv) “edema” (ED), v) “tumor core” (TC) as a combination of ET, NET, and CC, vi) the entire tumorous region, or the so-called “whole tumor” (WT).
@@
    Online evaluation platform (Synapse https://www.synapse.org/Synapse:syn53708249/wiki/626323)